# The Impact of Prompt Programming on Function-Level Code Generation

### v1

@@ V-A Correctness
A central question for assessing the value of prompt techniques is how likely a (combination of) techniques is to lead to code that (i) does not throw errors and (ii) passes the tests. To measure the correctness for different prompts, we use the well-established Pass@k metric [27]. This metric measures the likelihood that the LLM will generate a correct solution from the $k^{st}$ attempt. When $k=1$, then the metric becomes a measure of accuracy, as it calculates the number of passed functions divided by the total number of functions. In our study, we run the functions generated by the 32 combinations of prompt techniques on the tests provided by the CoderEval benchmark. Then, we collect the test results (pass or fail) and measure Pass@1 accordingly. Figure 5 shows Pass@1 results for all combinations of techniques (see Table I) for GPT-4o. Given that results between different models appear to be very consistent (see also Section V), we focus our discussion on one example model. However, results for the other models can be found in Appendix -A.
@@
It is evident from Figure 5 that the most important technique when it comes to correctness is the presence of a function signature. Combining the signature with other techniques, such as few-shot or chain-of-thought, is sometimes helpful to further increase the likelihood of a generated function being correct (albeit by a very small margin, e.g., adding chain-of-thought and package information to the signature only leads to an improvement of 0.4 percentage points). The best combination, with a Pass@1 of 57.9%, is the combination of signature and few-shot. We achieved the worst results in terms of correctness when using chain-of-thought alone, with a Pass@1 of 46.2%. It is surprising to note that the impact of prompt engineering techniques is overall lower than we would have expected — the difference between the best and worst combinations is merely 11.7 percentage points, i.e., prompt programming seems to have a noticeable impact in only a little over one in ten generation tasks.
Our findings also indicate that sometimes the addition of more information in the prompt leads to worse performance. For example, using only few-shot and signature performs better than if also packages are provided. Further, it is evident that techniques can interact in non-obvious ways. For example, both package information and CoT alone led to the worst Pass@1 results in the experiment. However, if these techniques are used in conjunction with a function signature, Pass@1 improves marginally over using only the signature in isolation.
Figure 6: Multi-linear regression results for the test results (pass or fail). Each point visualizes the coefficient estimate for the corresponding combination. The darker colors represent more conservative significance levels ($\alpha$). Zero-shot is not depicted, as it cannot be combined with other techniques.
To further investigate these interactions between factors in our experiment, we conducted a multi-linear regression analysis. Figure 6 shows the five prompt techniques in the study their interactions and their effect on the test result (pass or fail). For instance, “CoT:Persona” describes if the impact on test results comes from the interaction of CoT and Persona in a prompt, regardless of whether that prompt includes other prompt techniques. Similarly, “Sig.” (signature) refers to all prompts that include at least the signature (including, for example, P23, the combination of few-shot and signature), and is not limited to prompts that only specify the signature.
@@
The remaining main factors (packages, chain-of-thought, and persona) do not have a statistically significant impact on any of the three models. However, it is interesting to observe that the impact of a persona as well as chain-of-thought even trends to the negative (coefficient estimate smaller than 0).
Key Findings (RQ2.1): The presence of a signature or few-shot has the clearest positive impact on correctness. The other prompt techniques in the study do not have a statistically significant impact on correctness. However, in general, the difference between “good” and “bad” prompt techniques is surprisingly low. Adding additional information to a prompt sometimes leads to worse performance.
Digging deeper into what causes generated functions to fail, Figure 7 displays the percentages of errors types encountered for each combination of prompt techniques. We show the four most common error types (AssertionError, TypeError, AttributeError, and ImportError) using the GPT-4o model (see Appendix -C for the other models). For example, 51% of the failed functions of prompts with few-shot, packages, and signature throw an AssertionError.
Figure 7: The percentages of error types that we observed in failed functions generated by different combinations of prompt techniques (GPT-4o).
@@
Taking a closer look at the different error types, we observe that while AssertionErrors generally occur at a higher rate than the other error types across all prompt techniques, they are particularly more frequent in prompts that include the function signature. In contrast, the absence of the function signature often leads to TypeErrors, primarily because the LLM misjudges the expected number or order of positional arguments when generating functions.
ImportErrors were observed more frequently in prompts that did not employ few-shot prompting. Interestingly, in a subset of cases, ImportErrors occurred even when packages were explicitly specified. To investigate this, we manually inspected five random prompts where packages were specified but still resulted in ImportErrors. We found that when the prompt indicated the use of a package that is local or unfamiliar to the LLM, the LLM hallucinated and attempted to import non-existent functions from the specified packages.
@@
Key Findings (RQ2.1): Including the signature or few-shot examples in prompts generally reduces errors, particularly TypeErrors, AttributeErrors, and ImportErrors. Providing package information can naturally reduce ImportErrors but may cause hallucinated imports if unfamiliar to the LLM.
@@ V-B Similarity
Beyond correctness, we believe that another important question is how similar generated functions are to the human-written baseline. We use the CodeBLEU score [29] to measure how similar the generated function is to the baseline in terms of syntax, semantics, and logic structure. CodeBLEU is a composite score that integrates the scores of: N-gram match (BLEU [41]), a weighted BLEU, and the Abstract Syntax Tree (AST) node match, and the dataflow match as a semantic similarity measure with different weights (by default, 25% for each of the four scores). In our analysis, we remove the signature of the generated function and the ground truth before measuring the similarity to avoid any bias toward the signature prompt technique. Consequently, we can no longer measure the dataflow similarity as the functions are no longer parsable (i.e., identifying parameters and their dataflow in the function). Therefore, we set the weight for the dataflow match to 0, and 1/3 for the three other match scores.
@@
Figure 8: The average CodeBLEU scores for the functions generated by each combination (Llama3).
Figure 9: The coefficient estimates from the multi-linear regression of prompt technique combinations that significantly impact the CodeBLEU and its components n-gram match (BLEU), weighted n-gram match (BLEU-weighted), and AST match scores. The combinations of prompt techniques that do not have a significant impact on the CodeBLEU or any of its components using any model are omitted for brevity. A complete plot can be found in Appendix -D.
@@
In Figure 8, we see that the overall CodeBLEU score is low for all approaches (varying between 12.2% and 17.9% for Llama3) indicating that generated solutions are largely different than how humans have solved the same tasks. Results for the other models are in Appendix -D. We observe that using any prompt technique increases similarity (i.e., zero-shot has the lowest similarity to the baseline for all three models). Consistently with correctness, combinations that include a signature lead to higher similarity. This is unsurprising, given that a predefined signature restricts the solution space for the LLM (which can be desirable or unwanted depending on context). Combining more techniques indeed seems to generally increase similarity.
To better understand the impact of the prompt techniques and their interactions on the code similarity, we now perform a general multi-linear regression to see how the different prompt techniques and their interactions can impact the CodeBLEU and the three constituent dimensions we use (n-gram, weighted n-gram, and AST match scores).
N-gram and weighted N-gram matches can give us insights into how similar the generated function to the baseline on a lexical level, i.e., the set of words such as coding style, variable naming, and/or in-line documentation in the function. The AST match focuses on the syntactic structure and code constructs in the function such as nested statements independently of their naming (e.g., loops, conditionals, operations).
Figure 9 shows the results of our multi-linear regression. We see that, regardless of the test results, the presence of a signature or persona in a prompt can significantly increase the CodeBLEU score. Chain-of-thought (CoT) seems to also positively impact the CodeBLEU score, but only for Llama3. The impact of the signature was shared across all dimensions of CodeBLEU. The impact of the persona and CoT was only observed for the lexical similarity, which explains the slightly lower $p$-values compared to the signature. In an interesting case, few-shot seems to lower the lexical similarity but increase the syntactic similarity, which results in no overall impact on the CodeBLEU. The impact of few-shot suggests that while it generates functions that may use different variable naming and style (lower lexical similarity), the structure and logic are close to the baseline (higher syntactic similarity).
@@
Key Findings (RQ2.2): The signature, persona, and chain-of-thought increase the overall similarity of the function to the baseline (i.e., code written by humans). Few-shot increases the syntactic similarity (structure of the function) and decreases the lexical similarity (variable naming).
@@ V-C Quality
For code smells, we run Pylint on the generated functions for each prompt in CodePromptEval, using the code smell IDs defined by Pylint. Then, we group the code smells for prompts that share the same prompt techniques, and finally, we filter code smells that make up less than 5% of the total number of code smells for the group.
TABLE III: List of Code Smells in generated functions.
@@
Warning W0621 Redefines name from outer scope.
@@
Refactoring R1705 Unnecessary ”else” after ”return”.
@@
We find 12 code smells that fulfill these criteria for all LLMs (see Table III). Most identified code smells are warning and convention code smells, but there is also one error and two refactoring smells. From this list, we decided to remove C0304 as it is present in all generated functions across all LLMs and is mostly an artifact of our generation pipeline.
@@
In Figure 10, we show what percentage of functions have at least one instance of each code smell. For reasons of brevity, we focus on GPT-4o (Llama3 and Mistral’s in Appendix -B).
We note that 72% of the functions generated using the few-shot technique contain C0116 code smells, indicating that these functions lacked a descriptive docstring (in contrast to only 17% of functions generated by chain-of-thought combined with a persona and package information). In general, we observe that prompts that apply the few-shot and signature prompt techniques generate functions with more code smells and, more specifically, warning and error code smells compared to other prompts.
@@
Key Finding (RQ2.3): While using CoT, persona, or package information leads to fewer correct solutions, these techniques lead to higher-quality code in terms of code smells.
We now turn towards the cyclomatic and cognitive complexity and compare the complexity of generated solutions to the complexity of the human-written baseline. In Table IV, we show the $p$-values resulting from the paired Wilcoxon test to assess the statistical significance of differences between the cyclomatic complexity of the generated functions and ground truth. We only show the (combinations of) prompt techniques that had a significant impact on the cyclomatic complexity for at least two LLMs ($\alpha=0.05$).
TABLE IV: Cyclomatic complexity analysis using Wilcoxon test and A12 Vargha Delaney for the effect size (N- Negligible, S- Small, M- Medium, L- Large). $\downarrow$ indicates a reduction in complexity, $\emptyset$ indicates no statistical difference.
@@
Package 0.0012 0.403 $\downarrow$ (S) 0.0008 0.402 $\downarrow$ (S) 0.0407 0.417 $\downarrow$ (S)
CoT, Package 0.0366 0.446 $\downarrow$ (S) 0.0147 0.440 $\downarrow$ (S) 0.0211 0.410 $\downarrow$ (S)
Zero-shot 0.0143 0.433 $\downarrow$ (S) 0.0012 0.392 $\downarrow$ (S) 0.0402 0.407 $\downarrow$ (S)
CoT, Persona 0.0576 0.462 $\emptyset$ 0.0100 0.432 $\downarrow$ (S) 0.0062 0.433 $\downarrow$ (S)
Persona, Sig. 0.0999 0.466 $\emptyset$ 0.0358 0.436 $\downarrow$ (S) 0.0080 0.439 $\downarrow$ (S)
Persona 0.0857 0.464 $\emptyset$ 0.0086 0.411 $\downarrow$ (S) 0.0259 0.436 $\downarrow$ (S)
Sig. 0.1593 0.453 $\emptyset$ 0.0094 0.425 $\downarrow$ (S) 0.0157 0.434 $\downarrow$ (S)
Package, Sig. 0.0426 0.457 $\downarrow$ (N) 0.0156 0.417 $\downarrow$ (S) 0.0175 0.448 $\downarrow$ (S)
Persona, Package 0.0283 0.438 $\downarrow$ (S) 0.0039 0.404 $\downarrow$ (S) 0.0661 0.460 $\emptyset$
CoT 0.0191 0.447 $\downarrow$ (S) 0.0559 0.448 $\emptyset$ 0.0095 0.417 $\downarrow$ (S)
We use Vargha Delaney A12 measure [43] to understand the nature of the impact (reduces or increases complexity) and to quantify the effect size (Negligible ($A_{12}\geq 0.45$), Small ($0.36\leq A_{12}<0.45$), Medium ($0.36\leq A_{12}<0.29$), or Large ($A_{12}\leq 0.29$)). Vargha Delaney A12 is a probability measure (that was later adopted as an effect size measure), which describes the probability that one level (generated function complexity) is greater than a corresponding value in another level (ground truth complexity). If the A12 is less than 0.50, it means that the values of the first level are lower than the second level, and the lower the score is, the larger the effect size. This allows us to see if the prompts generate functions with a significantly lower or higher complexity as the ground truth, or with a comparable complexity when no significance is observed. Complete complexity results are in Appendix -E.
TABLE V: Cognitive complexity analysis using Wilcoxon test and A12 Vargha Delaney for the effect size (N- Negligible, S- Small, M- Medium, L- Large). $\downarrow$ indicates a reduction in complexity, $\emptyset$ indicates no statistical difference.
@@
Package, Sig. 0.0365 0.449 $\downarrow$ (S) 0.0001 0.382 $\downarrow$ (S) 0.0175 0.448 $\downarrow$ (S)
Few-shot, Persona 0.0364 0.442 $\downarrow$ (S) 0.0327 0.412 $\downarrow$ (S) 0.2305 0.447 $\emptyset$
CoT, Persona 0.3678 0.501 $\emptyset$ 0.0026 0.419 $\downarrow$ (S) 0.0062 0.433 $\downarrow$ (S)
CoT, Package 0.3300 0.485 $\emptyset$ 0.0256 0.444 $\downarrow$ (S) 0.0211 0.410 $\downarrow$ (S)
CoT 0.3609 0.496 $\emptyset$ 0.0449 0.445 $\downarrow$ (S) 0.0095 0.417 $\downarrow$ (S)
Persona, Sig. 0.0883 0.456 $\emptyset$ 0.0025 0.416 $\downarrow$ (S) 0.0079 0.439 $\downarrow$ (S)
Persona 0.1156 0.493 $\emptyset$ 0.0032 0.393 $\downarrow$ (S) 0.0259 0.436 $\downarrow$ (S)
Package 0.0777 0.446 $\emptyset$ 0.0002 0.371 $\downarrow$ (S) 0.0407 0.417 $\downarrow$ (S)
Sig. 0.5717 0.485 $\emptyset$ 0.0003 0.398 $\downarrow$ (S) 0.0157 0.434 $\downarrow$ (S)
Zero-shot 0.1708 0.486 $\emptyset$ 0.0016 0.382 $\downarrow$ (S) 0.0402 0.407 $\downarrow$ (S)
@@
For cognitive complexity (see Table VII), we observe larger differences among the LLMs than between combinations of prompt techniques in general. There was only one combination of prompt techniques that reduced the cognitive complexity across all three LLMs with a small size effect. GPT-4o seems to generate functions with no or negligible differences to the ground truth. Mistral can reduce the cognitive complexity with a small effect size when the prompt does not include few-shot and a persona, packages or CoT applied in the prompt. In contrast, there are no clear trends or patterns among the prompts in Llama3 rather most of the prompt techniques seem to reduce the cognitive complexity with a small effect size. We conclude that Llama3 appears to lead to simpler solutions than the other models, particularly GPT-4o.
@@
Key Findings (RQ2.3): There are noticeable differences among models with regard to the complexity of the code they produce. Llama3 appears to produce simpler solutions systematically. There were no cases of increased complexity — LLM solutions were comparably complex to human-written code, or simpler.

### v2

@@ V-A Correctness
A central question for assessing the value of prompt techniques is how likely a (combination of) techniques is to lead to a correct code, meaning that it is both (i) syntactically and operationally correct (valid and does not throw errors) and (ii) semantically correct (passes the tests). To measure the correctness of different prompts, we use the well-established Pass@k metric [28]. This metric measures the likelihood of drawing $k$ passing functions from the results of $n$ number of generations (or repetitions).
In our study, we run the functions generated by the 32 combinations of prompt techniques over three repetitions ($n=3$) on the tests provided by the CoderEval benchmark. Then, we collect the test results (pass or fail) and measure Pass@1 ($k=1$) accordingly. Figure 5 shows Pass@1 results for all combinations of techniques (see Table I) for GPT-4o. Given that results between different models appear to be very consistent (see also Section V), we focus our discussion on one example model. However, results for the other models can be found in the supplemental material.
@@
It is evident from Figure 5 that the most important technique when it comes to correctness is the presence of a function signature. Combining the signature with other techniques, such as few-shot or chain-of-thought, is sometimes helpful to further increase the likelihood of a generated function being correct (albeit by a very small margin, e.g., adding chain-of-thought and few-shot examples to the signature only leads to an improvement of 0.1 percentage points). The best combination, with a Pass@1 of 57.5%, is the combination of signature and few-shot. We achieved the worst results in terms of correctness when using chain-of-thought alone, with a Pass@1 of 47.1%. It is surprising to note that the impact of prompt engineering techniques is overall lower than we would have expected — the difference between the best and worst combinations is merely 10 percentage points, i.e., prompt programming seems to have a noticeable impact in only a little over one in ten generation tasks.
Our findings also indicate that sometimes the addition of more information in the prompt leads to worse performance. For example, using only few-shot and signature performs better than if all possible prompt techniques are used. Further, it is evident that techniques can interact in non-obvious ways. For example, both package information and CoT alone led to the worst Pass@1 results. However, if these techniques are used in conjunction with a function signature, Pass@1 improves marginally over using only the signature in isolation.
Figure 6: Results from our regression analysis for the pass@1 scores. Each point visualizes the coefficient estimate for the corresponding combination. The darker colors represent more conservative significance levels ($\alpha$). Zero-shot is not depicted, as it cannot be combined with other techniques.
To further investigate these interactions between factors in our experiment, we conducted a multi-linear regression analysis. Figure 6 shows the five prompt techniques in the study their interactions and their effect on the pass@1 score. For instance, “CoT:Persona” describes if the impact on test results comes from the interaction of CoT and Persona in a prompt, regardless of whether that prompt includes other prompt techniques. Similarly, “Sig.” (signature) refers to all prompts that include at least the signature (including, for example, P23, the combination of few-shot and signature), and is not limited to prompts that only specify the signature.
@@
The remaining main factors (packages, chain-of-thought, and persona) do not have a statistically significant impact on any of the three models.
[…]
skipabove=5pt, linecolor=blue, roundcorner=2pt, innerleftmargin=0.3cm,innerrightmargin=0.4cm, linewidth=0.5pt, footnoteinside=false,backgroundcolor=blue!5 {mdframed}Key Findings (RQ2.1): The presence of a signature or few-shot has the clearest positive impact on correctness. The other prompt techniques in the study do not have a statistically significant impact on correctness. However, in general, the difference between “good” and “bad” prompt techniques is surprisingly low. Adding additional information to a prompt sometimes leads to worse performance.
Digging deeper into what causes generated functions to fail, Figure 7 displays the percentages of errors encountered for each combination of prompt techniques. We show the four most common error types (AssertionError, TypeError, AttributeError, and ImportError) using GPT-4o (other models in the supplemental material). For example, 48.1% of the failed functions of prompts with few-shot, packages, and signature throw an AssertionError.
Figure 7: The percentages of error types that we observed in failed functions generated by different combinations of prompt techniques (GPT-4o) over three runs.
@@
Taking a closer look at the different error types, we observe that while AssertionErrors generally occur at a higher rate than the other error types across all prompt techniques, they are particularly more frequent in prompts that include the function signature, meaning that failed functions by prompts with a function signature are able to run but fail their tests due to a semantics-related error. In contrast, the absence of the function signature often leads to ImportErrors as well as TypeErrors that primarily occur because the LLM misjudges the expected number or order of positional arguments when generating functions.
Interestingly, in a subset of cases, ImportErrors occurred even when packages were explicitly specified. To investigate this, we manually inspected five random prompts where packages were specified but still resulted in ImportErrors. We found that when the prompt indicated the use of a package that is local or unfamiliar to the LLM, the LLM hallucinated and attempted to import non-existent functions from the specified packages.
@@
[…]
skipabove=5pt, linecolor=blue, roundcorner=2pt, innerleftmargin=0.3cm,innerrightmargin=0.4cm, linewidth=0.5pt, footnoteinside=false,backgroundcolor=blue!5 {mdframed}Key Findings (RQ2.1): Including the signature or few-shot examples in prompts generally reduces errors, particularly TypeErrors, AttributeErrors, and ImportErrors. Providing package information can naturally reduce ImportErrors but may cause hallucinated imports if unfamiliar to the LLM.
@@ V-B Similarity
Beyond correctness, we believe that another important question is how similar generated functions are to the human-written baseline. We use the CrystalBLEU score [43] to measure how similar the generated function is to the baseline in terms of the syntax and semantics of the function combined. CrystalBLEU is seen as a stricter improvement over the older CodeBLEU metric [30]. In our analysis, we remove the signature of the generated function and the ground truth before measuring the similarity to avoid any bias toward the signature prompt technique.
@@
Figure 8: The average CrystalBLEU scores for the functions generated by each combination (Llama3).
Figure 9: The coefficient estimates from the linear mixed-effects regression of prompt technique combinations that significantly impact the CrystalBLEU score.
In Figure 8, we see that the average CrystalBLEU score across three runs is low for all approaches (varying between 16.8% and 7.5% for Llama3) indicating that generated solutions are largely different than how humans have solved the same tasks. Results for the other models are in the supplemental material.
@@
We observe that using any prompt technique increases similarity (i.e., zero-shot has the lowest similarity to the baseline for all three models). Consistently with correctness, combinations that include a signature lead to higher similarity. This is unsurprising, given that a predefined signature restricts the solution space for the LLM (which can be desirable or unwanted depending on context). Combining more techniques indeed seems to generally increase similarity. We also observe that few-shot can decrease the similarity, achieving a score of 8.5%. However, combining it with signature, chain-of-thought or persona can improve the similarity to 10% and above.
To better understand the impact of the prompt techniques and their interactions on the code similarity, we now perform a linear mixed-effects regression analysis to see how the different prompt techniques and their interactions can impact the CrystalBLEU while accounting for the within-group variation (random effects) that arise from the three runs of each LLM.
Figure 9 shows our linear mixed-effects regression results. We see that, regardless of the test results, the presence of a signature or persona in a prompt can significantly increase the CrystalBLEU score. Chain-of-thought (CoT) seems to also positively impact the CrystalBLEU score for Llama3, but significantly lower it for GPT-4o and Mistral. We also observed that the interaction between certain prompt techniques can either reinforce or counteract the effects seen when the techniques are used individually. For example, while both Signature and CoT independently increase similarity in Llama-generated functions, using them together in a prompt can reverse that positive effect and significantly reduce similarity.
[…]
@@
skipabove=5pt, linecolor=blue, roundcorner=2pt, innerleftmargin=0.3cm,innerrightmargin=0.4cm, linewidth=0.5pt, footnoteinside=false,backgroundcolor=blue!5 {mdframed}Key Findings (RQ2.2): The signature and persona increase the overall similarity of the function to the baseline (i.e., code written by humans). Few-shot decreases the similarity unless combined with chain-of-thought.
@@ V-C Quality
For code smells, we run Pylint on the generated functions for each prompt in CodePromptEval, using the code smell IDs defined by Pylint. Then, we group the code smells for prompts that share the same prompt techniques, and finally, we select the top 15 code smells that were the most frequent across all prompt techniques.
TABLE IV: List of Code Smells in generated functions.
@@
Refactoring R1705 Unnecessary “else” after “return”.
@@
We find 11 code smells that fulfill these criteria for all LLMs (see Table IV). Most identified code smells are convention code smells, but there is also one error, two warnings, and two refactoring smells. From this list, we decided to remove C0304 as it is present in all generated functions across all LLMs and is mostly an artifact of our generation pipeline.
@@
In Figure 10, we show what percentage of functions have at least one instance of each code smell. For reasons of brevity, we focus on GPT-4o (Llama3 and Mistral’s in the supplemental material).
We note that 71% of the functions generated using the few-shot technique contain C0116 code smells, indicating that these functions lacked a descriptive docstring (in contrast to only 22% of functions generated by chain-of-thought combined with a persona and package information). In general, we observe that prompts that apply the few-shot and signature prompt techniques generate functions with more code smells and, more specifically, warning and error code smells compared to other prompts.
@@
[…]
skipabove=5pt, linecolor=blue, roundcorner=2pt, innerleftmargin=0.3cm,innerrightmargin=0.4cm, linewidth=0.5pt, footnoteinside=false,backgroundcolor=blue!5 {mdframed}Key Finding (RQ2.3): While using CoT, persona, or package information leads to fewer correct solutions, these techniques lead to higher-quality code in terms of code smells.
We now turn towards the cyclomatic and cognitive complexity and compare the complexity of generated solutions to the complexity of the human-written baseline. In Table V, we show the $p$-values resulting from the paired Wilcoxon test to assess the statistical significance of differences between the cyclomatic complexity of the generated functions and ground truth. We only show the prompt techniques that had a significant impact on the complexity for at least two LLMs ($\alpha=0.05$). Complete results are in the supplemental material.
TABLE V: Cyclomatic complexity analysis using Wilcoxon test and A12 Vargha Delaney for the effect size (N- Negligible, S- Small, M- Medium, L- Large). $\downarrow$ indicates a reduction in complexity, $\emptyset$ indicates no statistical difference.
@@
Package 0.0015 0.403 $\downarrow$ (S) 0.0017 0.410 $\downarrow$ (S) 0.0050 0.412 $\downarrow$ (S)
Persona, Package 0.0136 0.442 $\downarrow$ (S) 0.0033 0.412 $\downarrow$ (S) 0.0046 0.424 $\downarrow$ (S)
Zero-shot 0.0172 0.445 $\downarrow$ (S) 0.0009 0.401 $\downarrow$ (S) 0.0001 0.384 $\downarrow$ (S)
CoT, Package 0.0188 0.448 $\downarrow$ (S) 0.0118 0.442 $\downarrow$ (S) 0.0016 0.409 $\downarrow$ (S)
CoT, Persona, Package 0.0213 0.451 $\downarrow$ (N) 0.0015 0.431 $\downarrow$ (S) 0.0067 0.433 $\downarrow$ (S)
Persona, Sig. 0.0755 0.475 $\emptyset$ 0.0230 0.440 $\downarrow$ (S) 0.0041 0.442 $\downarrow$ (S)
Persona 0.0524 0.466 $\emptyset$ 0.0015 0.401 $\downarrow$ (S) 0.0004 0.408 $\downarrow$ (S)
Sig. 0.0909 0.460 $\emptyset$ 0.0093 0.430 $\downarrow$ (S) 0.0076 0.444 $\downarrow$ (S)
Package, Signature 0.0295 0.460 $\downarrow$ (N) 0.0398 0.436 $\downarrow$ (S) 0.0143 0.458 $\downarrow$ (N)
We use Vargha Delaney A12 measure [47] to understand the nature of the impact (reduces or increases complexity) and to quantify the effect size (Negligible ($A_{12}\geq 0.45$), Small ($0.450.36\leq A_{12}<$ 0.45), Medium ($0.360.29\leq A_{12}<$ 0.36), or Large ($A_{12}\leq 0.29$)). Vargha Delaney A12 is a probability measure (that was later adopted as an effect size measure), which describes the probability that one level (generated function complexity) is greater than a corresponding value in another level (ground truth complexity). If the A12 is less than 0.50, it means that the values of the first level are lower than the second level, and the lower the score is, the larger the effect size. This allows us to see if the prompts generate functions with a significantly lower or higher complexity as the ground truth, or with a comparable complexity when no significance is observed.
TABLE VI: Cognitive complexity analysis using Wilcoxon test and A12 Vargha Delaney for the effect size (N- Negligible, S- Small, M- Medium, L- Large). $\downarrow$ indicates a reduction in complexity, $\emptyset$ indicates no statistical difference.
@@
Few-shot, Persona 0.0239 0.437 $\downarrow$ (S) 0.0106 0.403 $\downarrow$ (S) 0.2629 0.461 $\emptyset$
Few-shot, Persona, Package 0.0326 0.445 $\downarrow$ (S) 0.0210 0.419 $\downarrow$ (S) 0.8239 0.505 $\emptyset$
Persona, Sig. 0.1846 0.473 $\emptyset$ 0.0024 0.417 $\downarrow$ (S) 0.0438 0.465 $\downarrow$ (N)
CoT, Persona 0.4080 0.524 $\emptyset$ 0.0212 0.441 $\downarrow$ (S) 0.0052 0.445 $\downarrow$ (S)
CoT, Package 0.1694 0.485 $\emptyset$ 0.0245 0.448 $\downarrow$ (S) 0.0334 0.435 $\downarrow$ (S)
CoT 0.3593 0.516 $\emptyset$ 0.0444 0.458 $\downarrow$ (S) 0.0093 0.432 $\downarrow$ (S)
Persona 0.2361 0.513 $\emptyset$ 0.0027 0.398 $\downarrow$ (S) 0.0155 0.442 $\downarrow$ (S)
Package, Sig. 0.1617 0.468 $\emptyset$ 0.0003 0.406 $\downarrow$ (S) 0.0437 0.448 $\downarrow$ (S)
Package 0.1694 0.485 $\emptyset$ 0.0006 0.383 $\downarrow$ (S) 0.0267 0.426 $\downarrow$ (S)
Zero-shot 0.2522 0.504 $\emptyset$ 0.0008 0.385 $\downarrow$ (S) 0.0298 0.415 $\downarrow$ (S)
@@
For cognitive complexity (see Table VI), we observe larger differences among the LLMs than between combinations of prompt techniques in general. There was no combination of prompt techniques that reduced the cognitive complexity across all three LLMs. GPT-4o seems to generate functions with no or small differences to the ground truth. Mistral can reduce the cognitive complexity with a small effect size when the prompt does not include few-shot and a persona, packages or CoT applied in the prompt. In contrast, there are no clear trends or patterns among the prompts in Llama3 rather most of the prompt techniques seem to reduce the cognitive complexity with a small effect size. We conclude that Llama3 appears to lead to simpler solutions than the other models, particularly GPT-4o.
@@
[…]
skipabove=5pt, linecolor=blue, roundcorner=2pt, innerleftmargin=0.3cm,innerrightmargin=0.4cm, linewidth=0.5pt, footnoteinside=false,backgroundcolor=blue!5 {mdframed}Key Findings (RQ2.3): There are noticeable differences among models with regard to the complexity of the code they produce. Llama3 appears to produce simpler solutions systematically. There were no cases of increased complexity — LLM solutions were comparably complex to human-written code, or simpler.